# Unsupervised learning in particle physics

### v1

@@ 4.2 Detector simulation
CaloGAN consists of two neural networks: a generator (G) and a discriminator (D). It follows the DCGAN (Deep Convolutional GAN) architecture, with the inclusion of LAGAN (Locally Aware Global Anomaly Network) [^9] The key idea in LAGANs is the integration of location information into the GAN architecture, enabling the generation of images with improved spatial coherence and context-awareness. It concatenates spatial coordinates or spatial embeddings with the input at various layers of both the generator and discriminator networks. Spatial coordinates could be absolute pixel positions or relative positions within the image. Spatial embeddings are learned representations of location information that can capture spatial relationships within the image. By incorporating location information, the generator can learn to produce images with better spatial coherence, such as smooth transitions between objects or accurate placement of features. guidelines, as shown in Figs. 16 and 17. The generator takes a random noise vector (latent space) as input and maps the latent input to three grayscale image outputs with different numbers of pixels, as per the three calorimeter layers. The mismatch of dimensionality and granularity among the three longitudinal segmentations of the detector, demands separate streams of operations with suitably sized kernels. The generator architecture therefore has multiple independent processing paths, each resembling a LAGAN architecture, with a trainable attention mechanism focusing on the relationships between these parallel LAGAN streams. The LAGAN submodules consist of a 2D convolutional unit, for spatial feature extraction followed by two locally-connected units which focus on specific regions within the data, leveraging the spatial features extracted by the previous layer. The attention mechanism allows dependence among layers, helping in utilizing information from the previous layers to the generation of the subsequent layer readout. The discriminator D accepts the three images as inputs, along with the energy of the incoming particle, $E$. The images are augmented with a sub-differentiable version of the sparsity percentage, and minibatch discrimination [^10] Minibatch discrimination is used in the discriminator of GAN to increase diversity in generated samples and improve training stability. It addresses the problem of mode collapse, where the generator produces a limited variety of samples that are easily classified as fake by the discriminator. is used on both the generated images as well as the augmented ones with the output sparsity. The inputs are mapped to a binary output that classifies showers into real and fake, and a continuous output that calculates the total energy deposited in the three layers, and then compares it with the requested energy E.
@@
The CaloGAN model has the potential to significantly reduce the computational cost of simulating particle showers in electromagnetic calorimeters. Once trained the simulator can be used at a cost of as little as 12 microseconds per event: a five-order-of-magnitude speedup compared to the full simulation.
The non-convex min-max loss function associated with GAN sometimes can not converge to an optimum leading to "mode collapse", with the GAN learning to generate only a subset of the data. Moreover, the ability of GANs to truly reproduce the underlying distribution of the data is not clear mathematically as GANs can only learn the likelihood implicitly. Other state-of-the-art GAN architectures, such as WGAN-GP [98][99] (the generator minimizes the Wasserstein distance in WGAN), BIB-AE [100], and DCTRGAN, [101] have been proposed to help with training stability and mode collapse. The loss function in WGAN, comparable to the JSD, Eq. (37), is defined as:
@@
The authors have used a combination of Masked Autoencoders for Distribution Estimation (MADE) and Neural Spline Flows to maximize expressive power (see Sec. Appendix C MADE and neural spline flow for more details). They have used the same calorimeter configuration as CaloGAN and trained two separate normalizing flows, a smaller one to learn the distribution of energies deposited in the three layers of the calorimeter ensuring energy conservation. The second larger one is used to learn the shower shapes in each layer. They present a detailed comparison of CaloFlow results with CaloGAN and Geant4. The comparison is on the average shower shapes images of the 100K events, nearest neighbour images (selecting elements of the Geant4 set at random and looking for their nearest neighbours in the set of generated CaloFlow/CaloGEN samples) and histograms of relevant physical quantities to better assess the quality of the generated shower shapes and voxel level information. CaloGAN in some regions of the phase space shows a sign of mode collapse, with the results from CaloFLOW very close to their GEANT4 counterpart.
@@
Normalizing flows as we saw, learns a series of invertible transformations that map a simple distribution (e.g., Gaussian) to the target data distribution. The model is trained to maximize the likelihood of the observed data under these transformations. Vanilla NFs have difficulty scaling to higher-dimensional datasets, especially when used for the study of ultra-fine calorimeters [103]. It has been observed that for two of the three datasets in the recent Fast Calorimeter Simulation Challenge 2022 [104], the NF training time is prohibitive and would require significant research and development [105, 106]. CaloFlow though $\sim 50$ times faster than GEANT4, are found to be considerably slower than the GAN-based alternative. The Masked Autoregressive Flow (MAF) architecture used in CaloFlow though fast in density estimation, is slower in sampling. The Inverse Autoregressive Flow (IAF) on the other hand leads to a fast sampling and a slow density estimation. Implementing IAF in CaloFlow is not straightforward as the high dimensionality collider data cannot be trained using the negative log-likelihood objective, due to memory and time limitations (see Ref. [107] for more details). The authors in Ref. [107] have used the Probability Density Distillation method, alternatively known as the teacher-student training for CaloFlow v2. The MAF (teacher) leads to a fast density estimation, by quickly mapping the target data points to the latent space, $x\rightarrow z$, while the IAF (student) trains on the output of the trained MAF model, leading the IAF inverse to the mapping of the latent space back to the target data space, $z\rightarrow x^{\prime}$. The teacher-student training leads to CaloFlow v2 preserving the same high fidelity of CaloFlow, but also significantly faster comparable to the previous version as shown in Fig. 19.
@@
Score-based Generative Models (CaloSCORE): An alternative class of generative models is proposed in Ref. [108] known as score-based models. These models operate by estimating the gradient of the data distribution’s log-likelihood w.r.t. the model parameters. The training of the score-based generative model uses a diffusion process that involves iteratively transforming a simple distribution to match the target data distribution. The diffusion process is modeled using a stochastic differential equation, which describes how the probability density evolves. The training process is divided into discrete time steps, where the model updates the samples by applying a transformation that introduces noise. This transformation gradually deforms the simple distribution towards the target distribution. At each time step, the model estimates the gradient of the log-likelihood of the data distribution w.r.t the model parameters. The score is used to update the model parameters through gradient-based optimization methods, such as stochastic gradient descent (SGD). Importantly, the transformations applied during the diffusion process are designed to be reversible. This ensures that, given a sample from the target distribution, it is possible to trace back the diffusion process to obtain a sample from the simple distribution. The objective during training is to make the transformed samples at each time step indistinguishable from samples drawn from the true data distribution. This is achieved by minimizing a suitable divergence or distance measure between the transformed distribution and the true data distribution. By iteratively applying the diffusion process and updating the model parameters based on the estimated gradients, the score-based generative model learns to generate realistic samples that match the true data distribution.
@@
A detailed evaluation of the quality of generated samples using different CaloScore SDE implementations and a WGAN-GP model for fast calorimeter simulation is presented. Various distributions are used to assess the agreement between the Geant4 and generated samples, focusing on parameters such as total energy deposited, calorimeter hits, energy deposition distribution, and angular distribution of the calorimeter shower. They obtain good agreement between the CaloScore implementations and Geant4 predictions, with some deviations observed in specific scenarios. The WGAN-GP model shows higher deviations compared to CaloScore. CaloScore includes a scalable and stable training schedule, based on the minimization of the convex score-matching loss, and exact likelihood estimation. The major challenge to be addressed in CaloScore is the generation time, currently requiring hundreds of model evaluations to solve the reverse SDE. The potential for accelerating the generation process is noted and taken care of in Ref. [109], where the CaloScore architecture is updated through the use of the attention layer which produces higher-quality samples. The additional use of progressive distillation leads to an iterative reduction in the number of time steps required during sampling. This technique further shows that single-shot diffusion models can be achieved for fast and high-fidelity simulation in collider physics. Overall, CaloScore offers a new generative paradigm with the potential for producing realistic calorimeter showers in collider physics while addressing some limitations of existing generative models.
@@
ATLAS calorimeter simulation using GANs: The experimentalists at LHC have also established that a fast calorimeter simulation is essential to produce the number of simulated events needed to match the High Luminosity LHC. ATLAS first provided the AtlFastII tool [110] based on the FastCaloSim package [111]. This tool provided a parametrized simulation of the particle energy response and the energy distribution in the ATLAS calorimeter system hence reducing the calorimeter simulation time to a few seconds per event. The parametric simulations simulate the energy of a particle shower as a single step based on an underlying parametrization instead of simulating how every particle propagates and interacts inside the calorimeter volume. This tool has difficulty in modelling boosted objects and jets of particles reconstructed with large-radius clustering algorithms. Moreover, some observables in electromagnetic showers are not well reproduced. ATLAS has also proposed FastCaloGAN [112], based on GANs that can perform a full calorimeter simulation for three particle types (electrons, photons, and pions) for a wide range of energies. The FastCaloGAN tool is based on the conditional WGAN-GP algorithm, doing the conditioning on the true momentum of the input samples since WGAN has been demonstrated to provide good performance and stability of the training. A good level of agreement was observed on reconstructed variables for both single-particle showers and jets with the most notable difference coming from the total energy of $e/\gamma$ showers. The $e/\gamma$ showers were well reproduced by FastCaloSimV2, therefore the ATLAS experiment designed the AtlFastIII tool [103], a combination of the above two. It uses the $e/\gamma$ and hadron showers at high and low energies simulated by FastCaloSimV2 and medium energy hadrons by FastCaloGAN and also provides a high-performance fast simulation. It does a smooth interpolation between the two tools and is designed to simulate particle showers to a level of precision such that no sizeable differences from the Geant4 simulation can be resolved by the reconstruction algorithms. Additionally, it requires only 20% as much CPU as Geant4 to simulate an event.
@@
GANs at highly granular detectors: The 3DGAN conditional model is used as an imaging tool for electromagnetic calorimeter (ECAL) simulation of the CLIC detector in Ref. [113]. It is based on three-dimensional convolutions to capture the whole shower development along the three spatial dimensions corresponding to a specific particle energy. The discriminator in this adversarial network apart from discriminating the real and fake samples also performs two additional tasks, estimating the incoming particle energy ($E_{p}$) and the total energy measured by the calorimeter ($E_{cal}$). The authors have successfully generated three-dimensional images of energy showers, according to different energy and incident angle values. A good agreement is found between their 3DGAN and GEANT4 output, for both low-level variables describing the calorimeter energy response, i.e. the single cell energy distributions, and more complicated features such as average profiles of the electromagnetic showers along the three calorimeter axes. Additional works have been done in the context of planned highly granular International Large Detectors (ILD) for future linear colliders, or other detectors of high granularity, see Refs. [114, 115, 116, 117].
@@ 5 Conclusions and outlook
Particle physics in the era of high-luminosity LHC and other data-intensive experiments have led us to think of new ways to analyse and interpret experimental data. This, in turn, has paved the way for the natural synergy between particle physics and modern ML techniques. Supervised learning has been widely used for classifying signals from backgrounds at the colliders. However, all supervised methods work on labelled data, which rely heavily on our ability to generate synthetic datasets for training. We can not use supervised techniques on real data which has no labels like synthetic data. People have approached this issue in different ways. For example, depending on how much a priori information one has on the real data, one can use unsupervised, weakly supervised, or semi-supervised techniques. Alternatively, one can also use ML techniques to determine how well the simulations agree with data and then try to improve the simulations. Currently, all these approaches are actively being studied and applied in particle physics (see, e.g., Ref. [120]).
@@
Unsupervised clustering methods identify patterns and group objects in high-dimensional data from particle detectors—we use them to reconstruct and classify jets. Anomaly detection algorithms can identify rare or unexpected events in the data that may indicate the presence of NP or detector malfunctions. AEs and VAEs are commonly used for anomaly detection. AE models require a better understanding of the learned (latent) space. Some studies look at this aspect and other challenges associated with AEs [121]. Many weak-supervision techniques—particularly the CWoLa-based ones—assume that the anomaly is localised in a subspace of the entire feature space. As shown in Ref. [48], there is an asymptotic guarantee of optimality when the classifier is trained on mixed samples of signal and backgrounds, provided the samples have different signal fractions. However, in practice, the performance of such classifiers depends on several factors, particularly the signal fractions in SRs. There are studies that look into the correlations between these methods [122] (in the context of resonant anomaly detection)—a combination of these methods looks the most promising for anomaly detection. Comparisons between fully and partially unsupervised methods show that they are complementary and can be used in combination with each other [37]. ATLAS [123] and CMS [43] have started exploring these strategies for NP searches. By combining unsupervised learning with domain-specific knowledge and expertise, physicists can leverage the strengths of both approaches to extract meaningful insights from experimental data. The use of generative modelling in simulating particle collisions and detector unfolding is rising. Density estimation-based models like normalising flows (that can efficiently map a simple distribution to a target one) are being used for detector simulation. There are also classifier-based tools like Omnifold, which learn the mapping from simulation to observed data and then invert the mapping to unfold detector effects.
@@
In this review, we mainly focused on the interesting ideas in a small subset of the burgeoning field of data-driven physics models. These techniques look promising and their complete scope is yet to be explored. With the advent of noisy intermediate-scale quantum computers, the efficiency of the data-driven methods may take a higher leap in the near future. [^11] Quantum algorithms for ML [124, 125, 126, 127, 128], partiularly anomaly detection [129, 130, 131], is already being explored. Moreover, beyond the colliders, data-driven models, or ML methods in general, can play crucial roles in other domains like neutrino physics, string theory, lattice QCD, etc. (see Ref. [132] and the references therein).

### v3

@@ IV.2 Detector simulation
CaloGAN consists of two neural networks: a generator (G) and a discriminator (D). It follows the DCGAN (Deep Convolutional GAN) architecture, with the inclusion of LAGAN (Locally Aware Global Anomaly Network) [^9] The key idea in LAGANs is the integration of location information into the GAN architecture, enabling the generation of images with improved spatial coherence and context-awareness. It concatenates spatial coordinates or spatial embeddings with the input at various layers of both the generator and discriminator networks. Spatial coordinates could be absolute pixel positions or relative positions within the image. Spatial embeddings are learned representations of location information that can capture spatial relationships within the image. By incorporating location information, the generator can learn to produce images with better spatial coherence, such as smooth transitions between objects or accurate placement of features. guidelines, as shown in Figs. 15 and 16. The generator takes a random noise vector (latent space) as input and maps the latent input to three grayscale image outputs with different numbers of pixels, as per the three calorimeter layers. The mismatch of dimensionality and granularity among the three longitudinal segmentations of the detector, demands separate streams of operations with suitably sized kernels. The generator architecture, therefore, has multiple independent processing paths, each resembling a LAGAN architecture, with a trainable attention mechanism focusing on the relationships between these parallel LAGAN streams. The LAGAN submodules consist of a 2D convolutional unit, for spatial feature extraction followed by two locally-connected units which focus on specific regions within the data, leveraging the spatial features extracted by the previous layer. The attention mechanism allows dependence among layers, helping in utilizing information from the previous layers to the generation of the subsequent layer readout. The discriminator D accepts the three images as inputs, along with the energy of the incoming particle, $E$. The images are augmented with a sub-differentiable version of the sparsity percentage, and minibatch discrimination [^10] Minibatch discrimination is used in the discriminator of GAN to increase diversity in generated samples and improve training stability. It addresses the problem of mode collapse, where the generator produces a limited variety of samples that are easily classified as fake by the discriminator. is used on both the generated images as well as the augmented ones with the output sparsity. The inputs are mapped to a binary output that classifies showers into real and fake and a continuous output that calculates the total energy deposited in the three layers and then compares it with the requested energy E.
@@
The CaloGAN model has the potential to significantly reduce the computational cost of simulating particle showers in electromagnetic calorimeters. Once trained, the simulator can be used at a cost of as little as 12 microseconds per event—a five-order-of-magnitude speedup compared to the full simulation.
The non-convex min-max loss function associated with GAN sometimes can not converge to an optimum leading to "mode collapse", with the GAN learning to generate only a subset of the data. Moreover, the ability of GANs to truly reproduce the underlying distribution of the data is not clear mathematically as GANs can only learn the likelihood implicitly. Other state-of-the-art GAN architectures, such as WGAN-GP [112][113] (the generator minimizes the Wasserstein distance in WGAN), Bounded Information Bottleneck Autoencoder (BIB-AE) [^11] The BIB-AE architecture unifies architectures from GANs and VAEs with the principles of Information Bottleneck method. The primary aim is to balance the compression and the retention of relevant information in a way that is beneficial for representation learning tasks. The Encoder Network maps the input calorimeter images to some latent representation, with the decoder generating the calorimeter images from the latent space. The information bottleneck here refers to the principle that the latent representation is optimized by the model while also maximizing the mutual information between the input and output. [114], and GANs using the Deep neural networks using the Classification for Tuning and Reweighting protocol (DCTRGAN) [^12] The key idea here is to leverage classification-based tuning and reweighting strategies to enhance the quality and stability of the generated outputs from GANs. Classification-based tuning uses a classifier network trained alongside the GAN to guide the tuning of the GAN components, ensuring that the generated data aligns with desired class distributions. The reweighting part consists of adjusting the importance of different data points or features during training to address class imbalances or to emphasize certain characteristics in the generated data. For example, if certain classes are underrepresented, their corresponding data points can be given higher weights. The loss function for both the generator and discriminator can be adjusted using these weights, ensuring that the GAN learns to generate balanced and representative data.[115] have been proposed to help with training stability and mode collapse. The loss function in WGAN, comparable to the JSD, Eq. (37), is defined as:
@@
The authors have used a combination of Masked Autoencoders for Distribution Estimation (MADE) and Neural Spline Flows to maximize expressive power (see Appendix Appendix C MADE and neural spline flow for more details). They have used the same calorimeter configuration as CaloGAN and trained two separate normalizing flows, a smaller one to learn the distribution of energies deposited in the three layers of the calorimeter, ensuring energy conservation and another one to learn the shower shapes in each layer. They present a detailed comparison of CaloFlow results with CaloGAN and Geant4. The comparison is on the average shower shapes images of the 100K events, nearest neighbour images (selecting elements of the Geant4 set at random and looking for their nearest neighbours in the set of generated CaloFlow/CaloGEN samples) and histograms of relevant physical quantities to better assess the quality of the generated shower shapes and voxel level information. CaloGAN in some regions of the phase space shows a sign of mode collapse, with the results from CaloFLOW very close to their GEANT4 counterpart.
@@
Normalizing flows as we saw, learns a series of invertible transformations that map a simple distribution (e.g., Gaussian) to the target data distribution. The model is trained to maximize the likelihood of the observed data under these transformations. Vanilla NFs have difficulty scaling to higher-dimensional datasets, especially when used for the study of ultra-fine calorimeters [117]. It has been observed that for two of the three datasets in the recent Fast Calorimeter Simulation Challenge 2022 [118], the NF training time is prohibitive and would require significant research and development [119, 120]. CaloFlow, though $\sim 50$ times faster than GEANT4, is found to be considerably slower than the GAN-based alternative. The Masked Autoregressive Flow (MAF) architecture used in CaloFlow though fast in density estimation, is slower in sampling. The Inverse Autoregressive Flow (IAF), on the other hand, leads to a fast sampling and a slow density estimation. Implementing IAF in CaloFlow is not straightforward as the high dimensionality collider data cannot be trained using the negative log-likelihood objective, due to memory and time limitations (see Ref. [121] for more details). The authors in Ref. [121] have used the Probability Density Distillation method, alternatively known as the teacher-student training for CaloFlow v2. The MAF (teacher) leads to a fast density estimation, by quickly mapping the target data points to the latent space, $x\rightarrow z$, while the IAF (student) trains on the output of the trained MAF model, leading the IAF inverse to the mapping of the latent space back to the target data space, $z\rightarrow x^{\prime}$. The teacher-student training leads to CaloFlow v2 preserving the same high fidelity of CaloFlow, but also significantly faster comparable to the previous version as shown in Fig. 18.
@@
Score-based Generative Models (CaloSCORE): An alternative class of generative models is proposed in Ref. [122] known as score-based models. These models operate by estimating the gradient of the data distribution’s log-likelihood w.r.t. the model parameters. The training of the score-based generative model uses a diffusion process that involves iteratively transforming a simple distribution to match the target data distribution. The diffusion process is modelled using a stochastic differential equation, which describes how the probability density evolves. The training process is divided into discrete time steps, where the model updates the samples by applying a transformation that introduces noise. This transformation gradually deforms the simple distribution towards the target distribution. At each time step, the model estimates the gradient of the log-likelihood of the data distribution w.r.t the model parameters. The score is used to update the model parameters through gradient-based optimization methods, such as stochastic gradient descent (SGD). Importantly, the transformations applied during the diffusion process are designed to be reversible. This ensures that, given a sample from the target distribution, it is possible to trace back the diffusion process to obtain a sample from the simple distribution. The objective during training is to make the transformed samples at each time step indistinguishable from samples drawn from the true data distribution. This is achieved by minimizing a suitable divergence or distance measure between the transformed distribution and the true data distribution. By iteratively applying the diffusion process and updating the model parameters based on the estimated gradients, the score-based generative model learns to generate realistic samples that match the true data distribution.
@@
A detailed evaluation of the quality of generated samples using different CaloScore SDE implementations and a WGAN-GP model for fast calorimeter simulation is presented. Various distributions are used to assess the agreement between the Geant4 and generated samples, focusing on parameters such as total energy deposited, calorimeter hits, energy deposition distribution, and angular distribution of the calorimeter shower. They obtain good agreement between the CaloScore implementations and Geant4 predictions, with some deviations observed in specific scenarios. The WGAN-GP model shows higher deviations compared to CaloScore. CaloScore includes a scalable and stable training schedule based on the minimization of the convex score-matching loss and exact likelihood estimation. The major challenge to be addressed in CaloScore is the generation time, currently requiring hundreds of model evaluations to solve the reverse SDE. The potential for accelerating the generation process is noted and taken care of in Ref. [123], where the CaloScore architecture is updated through the use of the attention layer which produces higher-quality samples. The additional use of progressive distillation leads to an iterative reduction in the number of time steps required during sampling. This technique further shows that single-shot diffusion models can be achieved for fast and high-fidelity simulation in collider physics. Overall, CaloScore offers a new generative paradigm with the potential for producing realistic calorimeter showers in collider physics while addressing some limitations of existing generative models.
@@
ATLAS calorimeter simulation using GANs: The experimentalists at LHC have also established that a fast calorimeter simulation is essential to produce the number of simulated events needed to match the High Luminosity LHC. ATLAS first provided the AtlFastII tool [124] based on the FastCaloSim package [125]. This tool provided a parametrized simulation of the particle energy response and the energy distribution in the ATLAS calorimeter system hence reducing the calorimeter simulation time to a few seconds per event. The parametric simulations simulate the energy of a particle shower as a single step based on an underlying parametrization instead of simulating how every particle propagates and interacts inside the calorimeter volume. This tool has difficulty in modelling boosted objects and jets of particles reconstructed with large-radius clustering algorithms. Moreover, some observables in electromagnetic showers are not well reproduced. ATLAS has also proposed FastCaloGAN [126], based on GANs that can perform a full calorimeter simulation for three particle types (electrons, photons, and pions) for a wide range of energies. The FastCaloGAN tool is based on the conditional WGAN-GP algorithm, doing the conditioning on the true momentum of the input samples since WGAN has been demonstrated to provide good performance and stability of the training. A good level of agreement was observed on reconstructed variables for both single-particle showers and jets with the most notable difference coming from the total energy of $e/\gamma$ showers. The $e/\gamma$ showers were well reproduced by FastCaloSimV2. Therefore, the ATLAS experiment designed the AtlFastIII tool [117], a combination of the above two. It uses the $e/\gamma$ and hadron showers at high and low energies simulated by FastCaloSimV2 and medium energy hadrons by FastCaloGAN and also provides a high-performance fast simulation. It does a smooth interpolation between the two tools and is designed to simulate particle showers to a level of precision such that no sizeable differences from the Geant4 simulation can be resolved by the reconstruction algorithms. Additionally, it requires only 20% as much CPU as Geant4 to simulate an event.
@@
GANs at highly granular detectors: The 3DGAN conditional model is used as an imaging tool for electromagnetic calorimeter (ECAL) simulation of the CLIC detector in Ref. [127]. It is based on three-dimensional convolutions to capture the whole shower development along the three spatial dimensions corresponding to a specific particle energy. The discriminator in this adversarial network, apart from discriminating the real and fake samples, also performs two additional tasks, estimating the incoming particle energy ($E_{p}$) and the total energy measured by the calorimeter ($E_{cal}$). The authors have successfully generated three-dimensional images of energy showers, according to different energy and incident angle values. A good agreement is found between their 3DGAN and GEANT4 output, for both low-level variables describing the calorimeter energy response, i.e., the single cell energy distributions, and more complicated features such as average profiles of the electromagnetic showers along the three calorimeter axes. Additional works have been done in the context of planned highly granular International Large Detectors (ILD) for future linear colliders, or other detectors of high granularity, see Refs. [128, 129, 130, 131].
@@ V Conclusions and outlook
Particle physics in the era of high-luminosity LHC and other data-intensive experiments has led us to think of new ways to analyse and interpret experimental data. This, in turn, has paved the way for the natural synergy between particle physics and modern ML techniques. Supervised learning has been widely used for classifying signals from backgrounds at the colliders. However, all supervised methods work on labelled data, which rely heavily on our ability to generate synthetic datasets for training. We can not use supervised techniques on real data which has no labels like synthetic data. People have approached this issue in different ways. For example, depending on how much a priori information one has on the real data, one can use unsupervised, weakly supervised, or semi-supervised techniques. Alternatively, one can also use ML techniques to determine how well the simulations agree with data and then try to improve the simulations. Currently, all these approaches are actively being studied and applied in particle physics (see, e.g., Refs. [134, 135]).
@@
Unsupervised clustering methods identify patterns and group objects in high-dimensional data from particle detectors—we use them to reconstruct and classify jets. Anomaly detection algorithms can identify rare or unexpected events in the data that may indicate the presence of NP or detector malfunctions. AEs and VAEs are commonly used for anomaly detection. A comparison of the performance of AEs to other shallow ML models for different anomaly detection tasks is presented in Ref. [136]. AE models require a better understanding of the learned (latent) space. Some studies look at this aspect and other challenges associated with AEs [137]. Many weak-supervision techniques—particularly the CWoLa-based ones—assume that the anomaly is localised in a subspace of the entire feature space. As shown in Ref. [51], there is an asymptotic guarantee of optimality when the classifier is trained on mixed samples of signal and backgrounds, provided the samples have different signal fractions. However, in practice, the performance of such classifiers depends on several factors, particularly the signal fractions in SRs. There are studies that look into the correlations between these methods [138] (in the context of resonant anomaly detection)—a combination of these methods looks the most promising for anomaly detection. Comparisons between fully and partially unsupervised methods show that they are complementary and can be used in combination with each other [40]. ATLAS [139] and CMS [46] have started exploring these strategies for NP searches. By combining unsupervised learning with domain-specific knowledge and expertise, physicists can leverage the strengths of both approaches to extract meaningful insights from experimental data. The use of generative modelling in simulating particle collisions and detector unfolding is rising. Density estimation-based models like normalising flows (that can efficiently map a simple distribution to a target one) are being used for detector simulation. There are also classifier-based tools like Omnifold, which learn the mapping from simulation to observed data and then invert the mapping to unfold detector effects.
@@
In this review, we mainly focused on the interesting ideas in a small subset of the burgeoning field of data-driven physics models. These techniques look promising and their complete scope is yet to be explored. With the advent of noisy intermediate-scale quantum computers, the efficiency of the data-driven methods may take a higher leap in the near future. [^13] Quantum algorithms for ML [140, 141, 142, 143, 144], partiularly in anomaly detection [145, 146, 147, 148], is already being explored. Moreover, beyond the colliders, data-driven models, or ML methods in general, can play crucial roles in other domains like neutrino physics, string theory, lattice QCD, etc. (see Ref. [149] and the references therein).